OpenStudy (anonymous):

A group of friends is addressing wedding invitations. Casey can address 25 envelopes in one hour, Anne can address 43 envelopes in two hours, and Paula can address 70 envelopes in three hours. About how long will it take them to address 400 envelopes? 23.00 hours 6.47 hours 5.73 hours 17.39 hours I've tried this problem several times but I can't seen to remember how to do it. Can someone help without giving me the answer?

OpenStudy (anonymous):

Casey 1 envelop=1/25 hr Anne 1 envelop=1/43 hr and Paula 1 envelop=1/70 hr For all one envelop=1/25+1/43+1/70 hr if all starts work at same time...n proceed for 400.